OpenStudy (anonymous):

what is the product of sqrt 2 (6b sqrt 16 -2b)

OpenStudy (anonymous):

no this. @Hero

hero (hero):

\[\sqrt{2}(6b \sqrt{16} - 2b)\]

hero (hero):

Okay, so distribute the \(\sqrt{2}\)

hero (hero):

Actually, it would probably be much easier for you to reduce \(\sqrt{16}\)

hero (hero):

First, before doing anything else, just reduce \(\sqrt{16}\)

hero (hero):

Actually, I'll just post the steps I came up with

hero (hero):

\[\sqrt{2}(6b \sqrt{16} - 2b)\] Reduce \(\sqrt{16} = 4\) to get: \[\sqrt{2}(6b \dot\ 4 - 2b)\] Multiply 6 x 4 to get 24: \[\sqrt{2}(24b - 2b)\] Subtract 24b - 2b to get 22b: \[\sqrt{2}(22b)\]

hero (hero):

And you can re-write that as \(22b\sqrt{2}\)

hero (hero):

Which is one of your answer choices

OpenStudy (anonymous):

thanks so muchhh @hero